OpenStudy (anonymous):

Name the algebraic property demonstrated in the example below: If x = 3 and 3 = y, then x = y. Commutative Property of Addition Associative Property of Addition Transitive Property Symmetric Property

OpenStudy (mathstudent55):

Commutative: a + b = b + a Associative: a + (b + c) = (a + b) + c Transitive: If a = b, and b = c, then a = c. Symmetric: If a = b, then b = a.